OpenStudy (anonymous):

Can someone please verify my answer? At time t minutes, an object's distance from a point is given by s(t) feet. Which of the following represents the instantaneous rate of change of the velocity of the object at time t?

OpenStudy (acacia):

http://www.1728.org/velocity.htm I find this website very helpful you should give it a try

OpenStudy (anonymous):

The choices: \[s''(t)\] \[\lim_{h \rightarrow 0} \frac{ s''(t+h) - s''(t) }{ h }\] \[\lim_{h \rightarrow 0}\frac{ s(t+h) - s(t) }{ h }\] \[\frac{ s'(t+h) - s'(t) }{ h }\] \[\frac{ s'(t+1) - s'(t) }{ 1 }\]

OpenStudy (anonymous):

At first I thought it was the second limit, but that proved to be wrong.

OpenStudy (anonymous):

Based on your website I'm thinking it is the first limit I wrote as an option.

OpenStudy (anonymous):

The instantaneous rate of change of velocity = \(\bf v'(t)=a(t)\). So basically it's the derivative of velocity. We know that velocity is the derivative of position hence: \(\bf v'(t)=a(t)=s''(t)\)

OpenStudy (anonymous):

@iamsammybear

OpenStudy (anonymous):

Thank you so much! So it would be the first limit I wrote above.

OpenStudy (anonymous):

\(\bf a(t)=acceleration \ with \ respect \ to \ time\)

OpenStudy (anonymous):

It would be the first choice, yes.